OpenStudy (anonymous):

If delta H= -70.0 kJ and delta S = -0.400 kJ the reaction is spontaneous below a certain temperature. Calculate the temperature in Kelvins.

OpenStudy (aaronq):

\(\Delta G=\Delta H-T\Delta S\) spontaneity implies that \(\Delta G<0\) set: \(\Delta G=\Delta H-T\Delta S=0\) plug your values in and solve for T

OpenStudy (anonymous):

So it would be 280 K?

OpenStudy (aaronq):

Thats not what i got, double-check your algebra

OpenStudy (anonymous):

I got 28?

OpenStudy (aaronq):

\(-70.0-(T*-0.4)=0\) \(-70.0+T*0.4=0\) \(0.4T=70\) \(T=\dfrac{70}{0.4}=175~K\)

OpenStudy (anonymous):

how did T become positive?

OpenStudy (anonymous):

I see!!

OpenStudy (aaronq):

The two negative signs cancelled out: -(-0.4)=0.4

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (aaronq):

no problem!